# Autonomous Vehicle Controllers From End-to-End Differentiable Simulation

### v1

@@ I Introduction
When training AV controllers, it is common to treat the environment as a black-box function that is only used to evolve the states and provide an interactive element to the data collection process. Different RL algorithms adopt different strategies for handling it.
Figure 1: End-to-end learning of controllers. Our framework utilizes a differentiable simulator to learn autonomous vehicle controllers from the corrections between the simulated new states and the target states.
By using replay buffers to store past transitions, value-based [20], model-based [30], or policy gradient [32] methods are oblivious to how the next state has been generated. They treat it as either a target or some state from which to bootstrap a value estimate. This approach is general, but makes training slow and sample-inefficient, since the world dynamics and a behavioural policy have to be learned from the data. Imitation learning [5], being essentially supervised learning on expert actions, is more sample-efficient but suffers from generalization difficulties and avoids learning virtually any kind of dynamics information.
If the environment dynamics are known one would likely be able to harvest the best aspects from each of these methods. In fact, in a differentiable environment one can optimize the policy directly using gradient descent, just via supervision from expert agent trajectories, as shown in Fig. 1. The benefits of this are: 1) Obtaining an explicit policy for continuous control, 2) Fast inference, since there is no planning at test time, 3) Unbounded policy precision, due to the dynamics not being approximated, 4) Facilitating more grounded learning by incorporating the dynamics directly into the training loop. This occurs by computing gradients of the dynamics, along with the usual derivatives, within the backpropagation, and using them to update the policy.
@@
Waymax [10] was recently introduced as a differentiable, large-scale autonomous driving simulator. In this work, we utilize its differentiability to train controllers using Analytic Policy Gradients (APG), reaping all the benefits mentioned above. We cleverly design our model to not require differentiating through the simulator state, only through the dynamics. Combining this with a recurrent architecture that propagates temporal information across long simulated trajectories allows us to flow the dynamics gradients to earlier steps (trace the blue arrows in Fig. 3), improving training speed and performance. To the best of our knowledge, we are the first to apply APG on such a large-scale task, obtaining strong competitive performance comparable to larger heavily-engineered models which do not make use of simulators.
@@ II Related Work
RL methods that consider the environment as a black box, have different ways to learn a policy without the gradients of the dynamics, as shown in Table I. Because of that, most of them do not incorporate any world priors that are otherwise contained in the simulator. Compared to them, the analytic policy gradients approach (APG) is stable, sample efficient, and fast at test time. It has been used for trajectory tracking in drones [36], and is increasingly being used in differentiable physics simulators for robotics [8, 9, 12, 22, 18].
@@
Model-based [30, 26, 39] Offline, planning Slow, bounded by precision of dynamics model
@@
Modern approaches to trajectory prediction. Recently, trajectory prediction for AVs has been dominated by transformer models [29, 38, 24, 13]. Specifically, MTR [29] uses a transformer encoder to predict an initial per-agent trajectory and collect map features alongside it. It then refines this trajectory using a transformer decoder. MVTA [34] adapts MTR for closed-loop simulation, as required by WOSAC [21], by introducing receding-horizon predictions and variable-length history aggregation. Other models focus on collision-avoidance [4] or interactivity between agents in order to improve performance [15]. All these methods directly predict future locations, not actions, and therefore do not rely on any simulations.
@@ III Method
[…]
@@ III-A Differentiating through Waymax
Obtaining gradients. We apply our APG method in the Waymax simulator [10]. Being implemented in Jax [3], it is relatively flexible in choosing which variable to differentiate and with respect to what. That being said, many of the obtainable derivatives are not meaningful. For example,
@@
    The derivatives of future agent locations with respect to current traffic light states or roadgraphs are all zero, because the simulator dynamics (e.g., bicycle or delta dynamics) do not depend on the roadgraph or the traffic lights. There are no rigid-body collisions.
@@
What is meaningful and useful is to take the gradients of future controllable agent locations with respect to their current actions, $\partial\mathbf{s}_{t}/\partial\mathbf{a}_{t-1}$. These are precisely the derivatives needed for objective (1) and hence we focus on them. We treat the dynamics as another function in the computation graph, while ensuring that the relevant code structures for training are registered in Jax’s pytree registry, so that any tracing, just-in-time compilation, or functional transformations like grad can be applied on them.
(a) Resetting based on distance.
(b) Resetting based on time.
Figure 4: Resetting agent’s state during incremental learning. We can do so either (a), based on distance, whenever the agent goes too far from the log trajectory, or (b) every few timesteps. A loss (shown in grey) is computed on all timesteps, but the collected trajectory (red) is discontinuous.
@@ III-C Model overview
Agent processing. For the multi-agent experiments described in Sec. IV-A we slightly adapt our architecture to allow training with $N$ agents but evaluating with $M$. We adopt a small transformer that forces each agent to attend to the other ones in parallel. However, at the last transformer block we use a fixed number of learned queries which attend over the variable number of agent keys and values, effectively soft-clustering them, and becoming independent of their number. This allows us to train with 32 agents but evaluate with 128, as required by [7].
@@ IV Experiments
APG. If we have access to a simulator we can perform roll-outs at train time and supervise the resulting trajectories $\{\textbf{s}_{t}\}_{t=1}^{T}$ with $\{\hat{\textbf{s}}_{t}\}_{t=1}^{T}$. This is sufficient to make the controller robust and allows it to reproduce the trajectory at test time on its own. If the policy is stochastic, the learning process becomes less sample-efficient because it becomes exponentially less likely that in a far away $t$, a given location $(x,y)$ is reached. Thus, as shown in Figure 5 converging to the GT trajectory does not happen uniformly from all directions, but rather sequentially. To improve sample-efficiency, as a form of curriculum, we snap back the simulated trajectory whenever it goes beyond a threshold $\xi$ with respect to $\{\hat{\mathbf{s}}_{t}\}_{t=1}^{T}$, as in Fig 4a. Since resetting the simulated trajectory to the GT one breaks the computation graphs, we detach the gradients as shown previously in Fig. 2.
@@ IV-A Large-scale experiments in Waymax
Experiment setup. The Waymax simulator uses the underlying Waymo Open Motion Dataset (WOMD) [7] to initialize the simulator state. It contains close to 500K training scenarios, each of length 9 seconds and a framerate of 10 Hz. The history length is 10 frames (1 second) and the goal is to predict the next 80 frames. To train the model, we run multiple simulations, one per scenario, in parallel (this being the simulated minibatch) and use Adam to optimize them. We iterate through all scenarios (the equivalent of an epoch) multiple times. At test time, we simulate multiple roll-outs and compute the average displacement error (ADE), offroad rate, and collision rate for each mode. A collision rate of 20% means that 20% of all trajectories have some objects colliding at some timestep in them. Since the future is uncertain and we want to make sure that the GT locations are covered by at least one of our modes, we report the eval metrics from the best mode. Thus, we evaluate on minADE, min offroad rate, and min collision rate.
@@
Results. Our main results are shown in Table II. We compare behavioural cloning (BC) with APG on two tasks – planning, where we learn a controller only for the self-driving car, and multi-agent (MA), where we learn a single controller for all agents. For the latter setting, we project the roadgraph and agent coordinates in each agent’s own coordinate frame. Hence, our model is agent-centric. Agents’ actions are selected independently, where each agent considers the location of all others. Multi-agency results from all of them being controlled by the policy.
Training and evaluation specifics. For the planning task we evaluate with and without IDM [10], which makes other agents reactive to the ego-vehicle, instead of simply following the log trajectories. For MA-BC the training loss is applied over all simulated agents which have a valid GT action. For MA-APG it is applied over all agent transitions for which both the current sim agent and log agent are considered valid by the simulator. The general principle is that we must supervise the transitions of all agents controlled by the policy. We evaluate MA in two settings – on all valid objects, and on all objects for which is_modeled is set, according to WOMD [21].
In general, our method, outperforms those based on behavioural cloning (BC), in most cases obtaining better ADE, collision and offroad rates. It is also more certain in terms of which actions to take, leading to lower variability between individual trajectory modes. Fig. 7 shows that for our APG method, the log-divergence and overlap rates are almost independent of the number of modes. BC displays an obvious dependence and requires many modes in order for one of them to be accurate. And while with multiple modes it produces lower log-divergence for the planning task, the overlap rates are still considerably worse, indicating that the trajectory is not expert-like. These results suggest that APG tends to produce more confident trajectories, which in turn allows us to save on computation, as we need to evaluate fewer modes to get a sense of the mean predicted future.
@@
Robustness to noise and trajectory length. We perform an ablation where we make the environment dynamics stochastic at test time by adding different levels of noise to the next agent locations. Fig. 6 shows that our APG method compares favourably to BC. For very small noise levels the performance is not affected at all. Additionally, in Table V we show that training on half of each sequence, instead of the full one, reduces performance only marginally.
@@
Qualitative study. We showcase samples from our method in Fig. 8 and observe that in the planning setting APG realizes accurate stochastic trajectories. Errors occur mostly from not predicting the correct acceleration, not the steering, which is prevalent in the BC models and causes the agent there to swerve. The MA-APG trajectories are relatively less accurate because, unlike in planning, when all agents are controlled, any prediction will have an effect on any agent, making training more difficult.
@@ V Conclusion
In this work, we leveraged the differentiable nature of Waymax to train AV policies for planning and multi-agent control. We proposed measures – incremental learning, and a specific architecture based on an RNN for the temporal information transfer, and a transformer, for the spatial mixing of agent features – to efficiently train long episodes in diverse settings. Results suggest that our model is accurate and robust, while at the same time being lightweight and fast at inference time.

### v2

@@ I Introduction
When training AV controllers, it is common to treat the environment as a black-box function that is only used to evolve the states and provide an interactive element to the data collection process. Different RL algorithms adopt different strategies for handling it. Specifically, value-based [21], model-based [31], or policy gradient [34] methods are oblivious to how the next state has been generated. They treat it as either a target, or some state from which to bootstrap a value estimate. This approach is general, but makes training slow and sample-inefficient, since the world dynamics and a behavioural policy have to be learned from the data. Imitation learning [5], being essentially supervised learning on expert actions, is more sample-efficient but still avoids utilizing the environment dynamics to help learn a policy.
Figure 1: End-to-end learning of controllers. Our framework uses the gradients of the dynamics in a differentiable simulator to learn vehicle controllers from the corrections between the simulated new states and the target states.
If the environment dynamics are known, one would likely be able to harvest the best aspects from each of these methods. In fact, in a differentiable environment one can optimize the policy directly using gradient descent, just via supervision from expert agent trajectories, as shown in Fig. 1. The benefits of this are: 1) Obtaining an explicit policy for continuous control, 2) Fast inference, since there is no planning at test time, 3) Unbounded policy precision, due to the dynamics not being approximated, 4) Facilitating more grounded learning by incorporating the dynamics directly into the training loop. This occurs from mixing the gradients of the dynamics with the usual derivatives, within the backpropagation, and using them to update the policy.
@@
Waymax [10] was recently introduced as a differentiable, data-driven, large-scale autonomous driving simulator. In this work, we utilize its differentiability to train controllers using Analytic Policy Gradients (APG), reaping all the benefits mentioned above. Our proposed architecture is recurrent and propagates temporal information across long simulated trajectories, which allows the dynamics gradients to flow back to earlier steps (trace the blue arrows in Fig. 3), improving training speed and performance. To the best of our knowledge, we are the first to apply APG on such a large-scale task, obtaining strong competitive performance comparable to larger heavily-engineered AV models which do not make use of simulators.
@@ II Related Work
APG. Most RL methods that consider the environment as a black box (shown in Table I) do not incorporate any world priors otherwise contained in the simulator. Compared to them, the analytic policy gradients approach (APG) is stable, sample efficient, and fast at test time. It has been used for trajectory tracking in drones [39], including from visual input [11], quadruped control, [32], and is increasingly being used in differentiable physics simulators for robotic manipulation and navigation [8, 9, 13, 23, 19].
@@
Model-based [31, 27] Offline, planning Slow, bounded by precision of dynamics model
@@
Current approaches to trajectory prediction. Recently, trajectory prediction for AVs has been dominated by transformer models [30, 41, 25, 14]. Specifically, MTR [30] uses a transformer encoder to predict an initial per-agent trajectory and collect map features alongside it. It then refines this trajectory using a transformer decoder. MVTA [37] adapts MTR for closed-loop simulation, as required by WOSAC [22], by introducing receding-horizon predictions and variable-length history aggregation. Other models focus on collision-avoidance [4] or interactivity between agents in order to improve performance [16]. All these methods directly predict future locations, not actions, and therefore do not rely on any simulations.
@@ III Method
$\text{ where }{\textbf{s}}_{t}$ $=\text{Env}\left({\textbf{s}}_{t-1},\mathbf{a}_{t-1}\right)\text{ and }\textbf{a}_{t-1}\sim\pi_{\theta}\left(\textbf{s}_{t-1}\right)$
@@ III-A Differentiating through Waymax
Obtaining gradients. We apply our APG method in the Waymax simulator [10]. Being implemented in Jax [3], it is relatively flexible in choosing which variable to differentiate and with respect to what. However, one is now faced with the task of choosing which gradients to use for training and how. Some are unavailable or not meaningful. For example,
@@
    The derivatives of future agent locations with respect to current traffic light states or roadgraph points are all zero, because the simulator dynamics (e.g., bicycle or delta dynamics) do not depend on the roadgraph or the traffic lights.
@@
What is meaningful and useful is to take the gradients of future controllable agent locations with respect to their current actions, $\partial\mathbf{s}_{t}/\partial\mathbf{a}_{t-1}$. These are precisely the derivatives needed for objective (1) and hence we focus on them. Since the simulator state is represented as a set of tensors, and the motion dynamics are computed as continuous functions over them, we treat the dynamics simply as another operation in the computation graph, and apply Jax’s tracing, just-in-time compilation, and functional transformations like grad over it without much effort.
Figure 4: Resetting agent’s state during incremental learning. Once every $n$ steps, we reset the agent’s position to the corresponding log state. Blue arrows show the GT trajectory, red is the discontinuous simulated trajectory, and gray lines show the supervision.
@@ III-C Model overview
Agent processing. For the multi-agent experiments described in Sec. IV-A we slightly adapt our architecture to allow training with $N$ agents but evaluating with $M$. We adopt a small transformer [36] that forces each agent to attend to the locations of the other agents in parallel. However, at the last transformer block we use a fixed number of learned queries which attend over the variable number of agent keys and values, effectively soft-clustering them, and becoming independent of their number. This allows us to train with 32 agents but evaluate with 128, as required by [7].
@@ IV Experiments
APG. If we have access to a simulator we can perform roll-outs at train time and supervise the resulting trajectories $\{\textbf{s}_{t}\}_{t=1}^{T}$ with $\{\hat{\textbf{s}}_{t}\}_{t=1}^{T}$. This is sufficient to make the controller robust and allows it to reproduce the trajectory at test time on its own. If the policy is stochastic, the learning process becomes less sample-efficient because it becomes exponentially less likely that in a far away $t$, a given location $(x,y)$ is reached. Thus, as shown in Figure 5 converging to the GT trajectory does not happen uniformly from all directions, but rather sequentially. To improve sample-efficiency, as a form of curriculum, we snap back the simulated trajectory whenever it goes beyond a threshold $\xi$ with respect to $\{\hat{\mathbf{s}}_{t}\}_{t=1}^{T}$. In Fig 4 the resetting is based on time (once every $n$ steps), but here we reset based on distance. We also detach the gradients as shown previously in Fig. 2.
@@ IV-A Large-scale experiments in Waymax
Experiment setup. The Waymax simulator uses the underlying Waymo Open Motion Dataset (WOMD) [7] to initialize the simulator state. It contains close to 500K training scenarios, each of length 9 seconds and a framerate of 10 Hz. The history length is 10 frames (1 second) and the goal is to predict the next 80 frames. To train the model, we run multiple simulations, one per scenario, in parallel (this being the simulated minibatch) and use Adam to optimize them. We iterate through all scenarios (the equivalent of an epoch) multiple times. At test time, we simulate multiple roll-outs and compute the average displacement error (ADE), offroad rate, and collision (also called overlap) rate for each mode. A collision rate of 20% means that 20% of all trajectories have some objects colliding at some timestep in them. Since the future is uncertain and we want to make sure that the GT locations are covered by at least one of our modes, we report the eval metrics from the best mode. Thus, we evaluate on minADE, min offroad rate, and min collision rate.
@@
Baseline. We highlight that to the best of our knowledge, we are the first to apply APG in the Waymax + WOMD setting [10, 7]. Hence, as the main baseline we set a non-APG method based on behavioural cloning – the currently dominant approach to train AV policies.
Evaluation settings. Our main results are shown in Table II. We compare behavioural cloning (BC) with APG on two tasks – planning, where we learn a controller only for the self-driving car, and multi-agent (MA), where we learn a single controller for all agents. For the latter setting, we project the roadgraph and agent coordinates in each agent’s own coordinate frame. Hence, our model is agent-centric. Agents’ actions are selected independently, where each agent considers the location of all others. Multi-agency results from all of them being controlled by the trained policy.
Training and evaluation specifics. For the planning task we evaluate with and without IDM [10], which makes other agents reactive to the ego-vehicle, instead of simply following the log trajectories. For MA-BC the training loss is applied over all simulated agents which have a valid GT action. For MA-APG it is applied over all agent transitions for which both the current sim agent and log agent are considered valid by the simulator. The general principle is that we must supervise the transitions of all agents controlled by the policy. We evaluate MA in two settings – on all valid objects, and on all objects for which is_modeled is set (the objects to predict for) as designated by WOMD [22].
Input features. The inputs to the model consist of the current agent locations, nearest roadgraph points, current traffic light states, and a heading direction to condition the agent roughly in terms of where it should go. The same conditioning is used in the BC baseline.
Results analysis. Our APG method outperforms those based on behavioural cloning (BC), obtaining better ADE, collision and offroad rates. In a few settings BC obtains a lower min ADE value, but qualitative results suggest that the agent swerves, producing unrealistic behaviour in terms of collision and offroad rates. The corresponding APG trajectories are considerably more realistic. Without IDM agents the testing setting is maximally similar to the training setting and APG performs best, with considerable improvement in overlap rate compared to BC. A design limitation of APG is that adding IDM during evaluation causes the learned motion dynamics to be different from the ones seen at test time, which naturally hurts performance. We view this as expected, given its supervised trajectory training and leave it as future work to perform additional finetuning with IDM agents.
APG is also more certain in terms of which actions to take, leading to lower variability between individual trajectory modes. Fig. 7 shows that APG requires many fewer modes in order to reach a given performance. Thus, by producing more confident trajectories, it allows us to save on computation.
@@
Robustness to noise and trajectory length. To assess potential sim-to-real transfer, we make the environment dynamics stochastic at test time by adding different levels of noise to the next agent locations. Fig. 6 shows that our APG method compares favourably to BC. For very small noise levels the performance is not affected at all. Additionally, in Table V we show that training on half of each sequence, instead of the full one, reduces performance only marginally.
@@
Qualitative study. We showcase samples from our method in Fig. 8 and observe that in the planning setting APG realizes accurate stochastic trajectories. Errors occur mostly from not predicting the correct acceleration, not the steering, which is prevalent in the BC models and causes the agent there to swerve. The MA-APG trajectories are relatively less accurate because, unlike in planning, when all agents are controlled, any prediction will have an effect on any agent, making training more difficult. Yet the APG performance is still better than the BC one, shown in Table IV.
@@ V Conclusion
In this work, we leveraged the differentiable nature of Waymax to train AV policies for planning and multi-agent control. We proposed measures – incremental learning, and a specific architecture based on an RNN for the temporal information transfer, and a transformer, for the spatial mixing of agent features – to efficiently train long episodes in diverse settings. Results suggest that our model is accurate and robust, while at the same time being lightweight and fast at inference time, enabling potential real-world usage.
Acknowledgements. This research was partially funded by the Ministry of Education and Science of Bulgaria (support for INSAIT, part of the Bulgarian National Roadmap for Research Infrastructure).